### v1

# Enhancing LLM Reasoning via Non-Human-Like Reasoning Path Preference Optimization

Abstract
Current approaches for strengthening LLM reasoning tend to introduce a training bias toward human-like reasoning trajectories. In step-wise preference optimization, in particular, dependence on human or higher-capacity model annotations for intermediate steps limits exploration of alternative, non-human-like reasoning paths and thus constrains achievable performance. Furthermore, through a small-scale pilot study, we observed that in approximately 75% of cases, the model’s first erroneous step occurs after the lowest-confidence point. This suggests that guiding the model at its lowest-confidence point before an error provides more accurate supervision than locating the first explicit error. In this paper, we propose Confidence-Guided Reasoning Path Preference Optimization (CGPO), a method that leverages a confidence signal to identify points of maximal uncertainty in the model’s reasoning process and applies self-generated, non-human-like reasoning-path guidance to mitigate trajectory drift. Our experiments span diverse models applied to both code and mathematical reasoning tasks. The results show that, with the same amount of training data, our method using data generated by a small model can achieve better performance in most cases compared with approaches using data generated by a strong model or human-annotated.
Enhancing LLM Reasoning via Non-Human-Like Reasoning Path Preference Optimization
Junjie Lu1thanks:  Equal contribution., Yuliang Liu2,411footnotemark: 1, Chaofeng Qu3, Wei Shen6, Zhouhan Lin2,5, Min Xu1thanks:  Corresponding author 1University of Technology Sydney, 2Shanghai Innovation Institute, 3Southeast University, 4Nanjing University, 5Shanghai Jiao Tong University, 6Independent Researcher {Junjie.Lu,Min.Xu}@uts.edu.au, liuyl03181@gmail.com
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across various domains, stemming from their ability to reason (Zelikman et al., 2022; Hao et al., 2023; Lewkowycz et al., 2022; Magister et al., 2022; Khalifa et al., 2025). To further enhance their reasoning capabilities, researchers have conducted extensive studies on both the training and inference phases of the models. Training methods typically focus on teaching the model to follow human instruction through supervised training (Yue et al., 2023; Luo et al., 2023), or on preference optimization of reasoning chains using reward models or human feedback, usually known as RLHF (Guo et al., 2025; Zhang et al., 2024a; Le et al., 2022). For inference methods, researchers often employ prompt tuning (Kojima et al., 2022; Imani et al., 2023), or intervene in the model’s generation process by assessing consistency (Wang et al., 2022; Chen et al., 2022; Taubenfeld et al., 2025a), predicting confidence (Weng et al., 2022; Zhu et al., 2025b), and leveraging expected outcomes (Madaan et al., 2023; Cheng et al., 2024).
While chain-of-thought (CoT) (Wei et al., 2023) has become the standard approach for addressing complex reasoning tasks, early training methods primarily focused on outcome optimization, with limited attention to the structure and quality of reasoning paths (Shojaee et al., 2025). One possible is that outcome-based approaches have yielded reasonably satisfactory performance on current benchmarks (Gao et al., 2024b; Hosseini et al., 2024), while another is that assigning fine-grained rewards remains costly. Nevertheless, there have been efforts to optimize finer-grained aspects of the reasoning process, like step-wise methods (Snell et al., 2024; Wang et al., 2024b) and token-wise methods (Zhang et al., 2024b; Lin et al., 2024).
Figure 1: The difference between human-like reasoning and non-human-like reasoning. Human-like methods emulate human cognitive processes and typically use predefined, readable anchors as delimiters to split the reasoning path. Nevertheless, it imposes substantial constraints on the content and format of data during sampling. In contrast, in this paper, non-human-like delineate reasoning steps are based on model confidence and provide reward supervision for tokens where the model is most uncertain, incurring lower cost and exhibiting broader applicability across various tasks.
Existing step-wise preference optimization approaches usually locate the model’s first error (Luo et al., 2024; Lu et al., 2024) and construct preference pairs accordingly, enabling focused and targeted optimization. However, these methods often involve substantial labeling overhead, posing a bottleneck in their practical applicability. Furthermore, our analysis of several incorrect answers in mathematical reasoning tasks reveals that models often exhibit confusion, shown by relatively low confidence in generation, well before about 75% of the actual errors occur. Examples are provided in Appendix A.1. This observation inspired us to develop a method capable of identifying issues earlier in a fully automated, human-free manner.
We propose Confidence-Guided Reasoning Path Preference Optimization (CGPO), which anchors preference optimization to model uncertainty rather than human alignment or process-error identification. Concretely, CGPO is grounded in two key principles: (1) enabling the model to identify and refine its own confusion, and (2) leveraging preference pair construction to guide optimization. By preference optimization, our approach introduces a thoroughly non-human-like reasoning path exploration, as shown in Figure 1.
Our experiments show that CGPO yields consistent gains on mathematical reasoning under the same sample counts. The MetaMath-llama model improves by 4.15% on GSM8K (Cobbe et al., 2021) and by 2.54% on MATH (Hendrycks et al., 2021) compared to the baseline. Importantly, our method uses the training data generated by the policy model itself, whereas the baseline uses the training data generated by stronger models or annotated by humans. Models trained with CGPO further demonstrate better generalization on Omni-Math. For the code generation task, we train the DeepSeek-Coder-Instruct-7B by using CGPO and observe a 2.1% improvement on LiveCodeBench (Jain et al., 2024) and 4.0% improvement on LeetCodeDataset (Xia et al., 2025). We also examine the scalability of our method by increasing the training data from 10k to 80k, which results in the data leading to continued improvements.
Our ablation studies and analyses further investigate the differences in effects between human-like and non-human-like splits. Moreover, our experiments reveal that, within a certain range, intermediate reasoning path-focused optimization leads to more substantial enhancement than optimizing the entire reasoning path.
Our main contributions are as follows:
  1. 1.
    We introduce CGPO, a method that automatically constructs preference data and performs reasoning path preference optimization for LLMs without relying on more powerful models or human supervision. CGPO can sample key non-human-like reasoning paths from the model to construct preference-learning data, therefore improving the model’s reasoning ability.
  2. 2.
    We demonstrate on several mathematical reasoning benchmarks that CGPO improves models’ reasoning ability. We also apply CGPO to code generation, showing that CGPO generalizes to broader domains where human-like reasoning paths are hard to define.
  3. 3.
    We build a simple and efficient training codebase for reproducing CGPO, together with multiple datasets of non-human-like reasoning paths across different models and the corresponding models trained on these data. We will release all code, datasets, and trained models to facilitate reproducibility and comparison upon acceptance of this paper.
Figure 2: CGPO overview. The core idea of our method fails in two parts: 1) explore the non-human-like reasoning path by determining the model reasoning step using the initial policy model $\pi_{0}$, and 2) construct chosen/rejected pairs to train $\pi_{0}$ with step-wise DPO training objective.
2 Related Works
Preference Optimization for LLM: Preference optimization methods have been investigated to improve LLM performance on complex tasks. PPO (Schulman et al., 2017) balances reward maximization with stability through a clipped objective, serving as a foundational method in RLHF but with considerable complexity. Subsequently, critic-free RL methods (Shao et al., 2024; Ahmadian et al., 2024; Hu et al., 2025) reduced complexity by using reward model outputs as immediate rewards for policy gradient updates. And methods like DPO (Rafailov et al., 2024) and SimPO (Meng et al., 2024) directly optimize the model on pairwise preference signals. The reduced computational overhead, combined with sufficiently strong performance, has enabled preference-based methods to achieve widespread success across a variety of domains like instruction following (Wu et al., 2024; Cheng et al., 2025), mathematical reasoning (Pang et al., 2024), and code generation (Zhang et al., 2025; Zhu et al., 2025a).
Confidence-Aware Methods: Model confidence, which is commonly used to quantify model uncertainty, is widely used in LLMs (Fu et al., 2025; Taubenfeld et al., 2025b; Kang et al., 2025; Li et al., 2025; Singhi et al., 2025). While confidence can be interpreted from multiple perspectives. Some approaches measure confidence directly using the predicted probability of the selected token (Farquhar et al., 2024; Liu et al., 2025), which is also the approach adopted in our work. Others rely on the model’s outputs. Xiong et al. (2024) proposed a systematic framework to express model confidence by generation, Yona et al. (2024) measures model confidence by consistency of answers. There are also several methods using the entropy of probability distribution or probe-based methods to measure model confidence in other tasks (Stolfo et al., 2024; Beigi et al., 2024).
Reasoning Trajectory Optimization: LLM reasoning is viewed as a foundational capability for improving performance on complex tasks, and numerous studies have investigated how to improve this ability (OpenAI et al., 2024; Yang et al., 2025): Supervised fine-tuning methods directly align model trajectory to annotated sequence (Yu et al., 2025a); RLHF methods typically force the model to optimize the reasoning trajectory towards a higher score (Yu et al., 2025b); and test time methods intervene in the generation process to optimize reasoning trajectories (Muennighoff et al., 2025). The aforementioned training approaches, especially RLHF, are generally concerned with outcome optimization and do not explicitly enforce correctness along the reasoning trajectories. Lai et al. (2024) directly optimizes an independent reasoning path, the nearest one of our method, while with a significant annotation challenge.
3 Methods
In this section, we first introduce the basic formulation, then describe the construction of the CGPO
training data, as illustrated in Figure 2. Finally, we introduce the training objective of CGPO
3.1 Preliminaries
Given an input problem $\mathbf{x}$, an initial policy model $\pi_{0}$ can generate a reasoning sequence under the temperature-based sample strategy. We have $\mathbf{y}=(y_{1},y_{2},\dots,y_{T}),\quad y_{t}\in\mathcal{V}$, where $\mathcal{V}$ denotes the vocabulary. Notably, the sampled token in the generation may not be the token with the top-1 probability. So that at decoding time step $t$, we define the model confidence $c_{t}$ in decoding by the sampled token $y_{t}$, as
$c_{t}\triangleq p(y_{t}|\pi_{0},\mathbf{x},y_{<t})$
3.2 Preference Pairs Construction
Step Definition. To construct non-human-like step-wise chosen/rejected preference pairs, we first need to clarify the definition of a step in our method. Given $\mathbf{x}$ and the corresponding generated $\mathbf{y}$, we can get the confidence sequence for $\mathbf{y}$, denoted by $\mathbf{c}=(c_{1},c_{2},\dots,c_{T})$. A confidence threshold is employed to split the reasoning steps, with $\tau$ determined based on the distribution of all confidence values $c$ across the dataset. The tokens with confidence below $\tau$ will be the split tokens. Then we get $\mathbf{y}=\mathbf{s}=(s_{1},s_{2},\dots,s_{J})$, where $s_{j}$ denotes a reasoning step that consists of several tokens.
Initial Reasoning Trajectory Determination. To align with the pair-wise preference optimization paradigm, based on the step definition, we choose the sequence before the most uncertain step of $\mathbf{s}$ to serve as the shared initial reasoning trajectory $s_{init}$ of the chosen/rejected pairs.
Chosen/Rejected Pairs Construction. We introduce a reward model $R$ to select the first token of the chosen/rejected pairs given $(\mathbf{x},s_{init})$. The Top-k candidate tokens, ordered by their logits, are denoted as $(v_{1},v_{2},\dots,v_{k})$. The reward for each token is obtained as $R(\mathbf{x},s_{init},v_{k})$. Then, we select the token with the highest score and the one with the lowest score as the starting tokens of the chosen/rejected pairs, denoted by $y^{+}$ and $y^{-}$ respectively.
Subsequently, $\pi_{0}$ will sample the following tokens of $(\mathbf{x},s_{init},y^{+})$ and $(\mathbf{x},s_{init},y^{-})$ before end-of-sequence tag <EOS> or the first token with confidence below $\tau$.
Finally, for each given input, we have two reasoning trajectories: the chosen one $(s_{init},s^{+},\dots)$ and the rejected one $(s_{init},s^{-},\dots)$.
3.3 Training Objective
Our goal is to optimize $\pi_{\theta}$ based on the triplets $(s_{init},s^{+},s^{-})$. Let $\mathcal{D}$ denote the collection of such triples. With a fixed reference model $\pi_{\text{ref}}$ and temperature $\beta>0$, we optimize $\pi_{\theta}$ using the DPO-like objective
$\mathcal{L}_{\text{CGPO}}(\theta)$ $=-\,\mathbb{E}_{(s_{init},s^{+},s^{-})\sim\mathcal{D}}\Big[\log\sigma\big(\beta(\Delta)\big)\Big],$
$\sigma(a)$ $=\frac{1}{1+e^{-a}}.$
where $\beta$ is the inverse temperature and
$\Delta$ $=\Delta_{\theta}-\Delta_{\text{ref}},$
$\Delta_{\theta}$ $\triangleq\log\pi_{\theta}(s^{+}\!\mid s_{init})-\log\pi_{\theta}(s^{-}\!\mid s_{init}),$
$\Delta_{\text{ref}}$ $\triangleq\log\pi_{\text{ref}}(s^{+}\!\mid s_{init})-\log\pi_{\text{ref}}(s^{-}\!\mid s_{init}).$
Sequence log-likelihoods use natural logs and decompose autoregressively as
$\log\pi(s^{\pm}\!\mid s_{init})\ =\ \sum_{t}^{seq}\log\pi\!\big(y_{t}\mid x,s_{init}),$
where the sum starts at the branching index $y^{\pm}$. This objective increases the likelihood of the R{R}-selected high-quality trajectory $s^{+}$ while decreasing that of $s^{-}$; the reference term $\Delta_{\text{ref}}$ stabilizes the relative preference signal. We use Algorithm 1 to demonstrate the entire pipeline of CGPO.
Algorithm 1 CGPO Pipeline
1:Dataset $\mathcal{X}$, $\pi_{0}$, $R$, temp, $\tau$, $k$, $\beta$
2:Trained policy $\pi_{\theta}$
3:Initialize $\pi_{\theta}\leftarrow\pi_{0}$, preference dataset $\mathcal{D}\leftarrow\emptyset$
4:for each input $\mathbf{x}\in\mathcal{X}$ do
5:  Sample $\mathbf{y}$ from $\pi_{0}$; record confidences $c_{t}=\pi_{0}(y_{t}\mid\mathbf{x},\mathbf{y}_{<t})$
6:  Split $\mathbf{y}$ into steps $\mathbf{s}=(s_{1},\dots,s_{J})$ using $\tau$
7:  Find the most uncertain step; set $s_{\text{init}}=\text{concat}(s_{1},\dots,y_{min_{c}})$
8:  Obtain Top-$k$ next-token candidates $(v_{1},\dots,v_{k})$
9:  Score candidates $r_{i}\leftarrow R(\mathbf{x},s_{\text{init}},v_{i})$; choose $y^{+}=\arg\max_{i}r_{i}$ and $y^{-}=\arg\min_{i}r_{i}$
10:  Expand the chosen branch: sample $^{+}, ss^{-}$ from $\pi_{0}$ starting at $^{+}, yy^{-}$, stopping at EOS or the first token with confidence $\tau$
11:  Add $(\mathbf{x},s_{\text{init}},s^{+},s^{-})$ to $\mathcal{D}$
12:end for
13:for $d=1$ to $D$ do
14:  Update $\pi_{\theta}$ by Step-DPO loss, $Loss(\mathbf{x},y^{+}_{d},y^{-}_{d})$
15:end for
16:return $\pi_{\theta}$
4 Experiments and Analysis
In this section, we first introduce our experimental setup, including datasets, models, baselines, metrics, and parameter settings. We then present the experimental results, followed by an analysis of the CGPO transferability, generalization, and scalability.
4.1 Experiments Setup
Datasets and Models: For mathematical reasoning tasks, we use the prompts from the Step-DPO-10k [^1] https://huggingface.co/datasets/xinlai/Math-Step-DPO-10K dataset as model inputs $\mathbf{x}$ to construct the preference dataset for CGPO training. MetaMath-Mistral-7B (Yu et al., 2024), MetaMath-LLaMA-8B (Liu et al., 2025), Qwen2-7B-SFT [^2] https://huggingface.co/xinlai/Qwen2-7B-SFT, Qwen1.5-32B-SFT [^3] https://huggingface.co/xinlai/Qwen1.5-32B-SFT and DeepSeekMath-Base-SFT [^4] https://huggingface.co/xinlai/DeepSeekMath-Base-SFT are employed as the policy models $\pi_{0}$. To ensure that the capabilities of the Reward Model (RM) align with the policy model and can provide token-level fine-grained reward, we adopt ASPRM-L [^5] https://huggingface.co/Lux0926/ASPRM-L as $R$ to select candidate tokens after $s_{init}$, preventing performance gains from an overly strong RM.
For code generation tasks, we use the prompts from the training set of the LeetCodeDataset as model inputs $\mathbf{x}$ to construct the CGPO preference dataset. The policy model $\pi_{0}$ is Deepseek-Coder-7B-Instruct-v1.5 (Guo et al., 2024) and we adopt ASPRM-D [^6] https://huggingface.co/Lux0926/ASPRM-D as our reward model to select candidate tokens. The trained models are evaluated on the test sets of LiveCodeBench [^7] https://huggingface.co/livecodebench and LeetCodeDataset [^8] https://huggingface.co/datasets/newfacade/LeetCodeDataset.
Baselines and Metrics: For the mathematics domain, two baselines are considered: the base model and the model trained using the Step-DPO method. For the code domain, only the base model serves as the baseline.
For all tasks, model performance is reported using the accuracy metric. In mathematical reasoning tasks, a prediction is considered correct only if model’s final answer exactly matches the ground truth. For code-generation tasks, the generated code is executed in a sandbox and deemed correct if it passes all test cases.
Parameter Settings: For each base model in the mathematical domain, we construct the training data as follows. All 10,795 prompts from Step-DPO-10k are fed through the model to generate one output per prompt, yielding 10,795 outputs. For each output, we locate the lowest confidence point and use the reward model to score the top 8 token candidates, then branch into two reasoning paths. Decoding proceeds until the confidence falls below the 2% of the model’s confidence distribution or <EOS> token is reached. This produces a preference dataset of 10,795 pairs for each base model, equivalent in size to Step-DPO-10k. For the code domain base model, we sample 4 times per prompt from the 2,641 LeetCodeDataset training examples and apply the same token selection and reward-model procedure to generate 10,564 preference pairs.
For MetaMath-Mistral-7B and MetaMath-Mistral-8B, we follow the training configuration released by Step-DPO, set $\beta$ to 0.4, use a learning rate of 5e-7, set the global batch size to 128, and train for 8 epochs to reproduce the Step-DPO method. For all other Step-DPO baselines, we use the publicly released model checkpoints by Step-DPO. For all CGPO models, we train for 4 epochs using a global batch size of 128 and a learning rate of 5e-7. For MetaMath-Mistral-7B, Qwen2-7B-SFT and Deepseek-Coder-7B-Instruct-v1.5, we set $\beta$ to 0.4. For other models, we set $\beta$ to 0.3. We use the AdamW optimizer and a cosine learning rate schedule, with a warmup ratio of 0.1.
Table 1: Performance comparison of various models on mathematical and code benchmarks. Red and Green numbers represent the performance improvement or decline compared to the base model.
Dataset Model Base + Step-DPO + CGPO
GSM8K MetaMath-Mistral-7B 77.1 76.7 (-0.4) 79.4 (+2.3)
MetaMath-Llama-8B 81.8 84.8 (+3.0) 85.3 (+3.5)
Qwen2-7B-SFT 87.9 88.4 (+0.5) 88.6 (+0.7)
Qwen1.5-32B-SFT 90.0 90.3 (+0.3) 90.4 (+0.4)
DeepSeekMath-Base-7B-SFT 86.3 86.5 (+0.2) 87.1 (+0.8)
MATH MetaMath-Mistral-7B 26.6 27.3 (+0.7) 27.0 (+0.4)
MetaMath-Llama-8B 39.3 40.2 (+0.9) 40.3 (+1.0)
Qwen2-7B-SFT 54.2 54.6 (+0.4) 54.8 (+0.6)
Qwen1.5-32B-SFT 54.2 55.7 (+1.5) 55.2 (+1.0)
DeepSeekMath-Base-7B-SFT 51.4 52.5 (+1.1) 51.8 (+0.4)
LiveCodeBench Deepseek-Coder-7B-Instruct-v1.5 19.3 / 19.7 (+0.4)
LeetCodeDataset Deepseek-Coder-7B-Instruct-v1.5 12.7 / 13.2 (+0.5)
4.2 Overall Results
The main results are shown in Table 1. For the mathematics reasoning task, our method outperforms Step-DPO on GSM8K and achieves the greatest improvement on MetaMath-Llama-8B, with a 4.3% increase compared to the base model. We attribute this phenomenon to the reward model and the policy model being trained on the same series of foundational models. Notably, when training MetaMath-Mistral-7B using the Step-DPO method, the resulting model’s performance actually declined compared to the base model. This may be because it is difficult for the weaker model to follow the human-like reasoning paths and learn from them in a preference way. However, our method still leads to an improvement compared to the base model, which demonstrates the advantage of the model learning from non-human-like reasoning paths through self-exploration. On the MATH dataset, our method outperforms Step-DPO on MetaMath-Llama-8B and Qwen2-7B-SFT, but performs less well on the other three models. That may be because the MATH dataset is more challenging than the GSM8K dataset. The 10k data volume is insufficient for fully exploring non-human-like reasoning paths, limiting the model improvement. Additional analysis experiments are conducted below to verify this.
Additionally, for the code generation task, the Deepseek-Coder-7B-Instruct-v1.5 trained with our method achieved 2.1% and 4.0% over the base model on the LiveCodeBench and LeetCodeDataset respectively. This demonstrates that, beyond mathematical reasoning tasks, CGPO can also enhance model capabilities in tasks such as code generation, where human-like reasoning paths are not easily defined, by enabling the model to explore non-human-like reasoning paths.
Table 2: Scalability on CGPO training data: models were trained with 10k, 20k, 40k, and 80k data samples, respectively.
Model Dataset Data Scale
10k 20k 40k 80k
MetaMath-Llama-8B GSM8K 85.3 85.3 85.6 86.4
MATH 40.3 40.8 41.2 41.2
Qwen2-7B-SFT GSM8K 88.6 88.7 88.9 89.5
MATH 54.8 54.9 55.1 55.3
4.3 Analysis of Scalability
To explore model performance at different exploration degrees, we sampled 2, 4, and 8 times for each prompt on the Step-DPO-10k dataset using MetaMath-Llama-8B and Qwen2-7B-SFT, thereby constructing training datasets of 20k, 40k, and 80k.
Table 2 presents model performance on the GSM8K and MATH datasets. We find that increasing the exploration degree gradually improves performance on both datasets, with a notable jump on GSM8K when the data volume increases from 40K to 80K. These results indicate that CGPO exhibits effective scalability.
Table 3: Performance comparison on out-of-domain dataset (Omni-Math).
Dataset Model Base + Step-DPO + CGPO
Omni-Math MetaMath-Mistral-7B 8.4 7.7 (-0.7) 8.8 (+0.4)
MetaMath-Llama-8B 11.4 11.1 (-0.3) 11.7 (+0.3)
Qwen2-7B-SFT 14.3 14.0 (-0.3) 14.6 (+0.3)
DeepSeekMath-Base-7B-SFT 13.7 14.7 (+1.0) 14.2 (+0.5)
Qwen1.5-32B-SFT 14.5 15.1 (+0.6) 15.2 (+0.7)
4.4 Analysis of Out of Datasets
To analyze the out-of-distribution generalization performance of CGPO and its performance on challenging mathematical tasks, we use the Omni-Math dataset (Gao et al., 2024a) is used to evaluate the models. Omni-Math is an evaluation benchmark that encompasses 4,428 Olympic-level competition problems of 10 different difficulty levels and 33 subfields.
As shown in Table 3, compared with Step-DPO, CGPO performs slightly lower only on DeepSeekMath-Base-7B-SFT, while outperforms it on all other models. This shows that CGPO exhibits superior out-of-distribution generalization and can effectively enhance the performance of weaker models on challenging benchmarks, even without strong model annotations.
Table 4: Results of training datasets constructed by different reward models: ASPRM and Math-Shepherd denote the models trained with preference data guided by these two reward models.
Model GSM8K MATH
MetaMath-Llama-8B   + ASPRM 85.3 40.3
+ Math-Shepherd 84.1 40.2
Qwen2-7B-SFT   + ASPRM 88.6 54.8
+ Math-Shepherd 88.2 54.7
4.5 Analysis of Reward Model
To explore the impact of different reward models on CGPO, we adopt the process reward model proposed by Math-Shepherd (Wang et al., 2024a) for comparison with ASPRM. The comparison results are presented in Table 4.
When Math-Shepherd is used instead of ASPRM, the performance of CGPO-trained model declines, although it remains better than the base model. Even with a weaker PRM, CGPO still shows an improvement over the base model. This improvement stems from precise token-level rewards, underscoring the significance of fine-grained evaluation.
Figure 3: The relationship between the confidence threshold and the trained model performance.
Figure 4: The relationship between the confidence threshold and the average number of tokens per pair.
4.6 Analysis of Stopping Threshold
We report the performance differences of the model under different confidence thresholds at which the sampling stops in Figure 3. For MetaMath-Llama-8B, performance keeps improving as the threshold increases. For Qwen2-7B-SFT, overall performance also shows an upward trend as the threshold rises, but the peak is achieved at a confidence threshold of 8%.
When we increase the confidence threshold, more samples will stop in the intermediate process rather than the generation of the <EOS> tag. This increases the proportion of non-human-like data, thereby enhancing the model’s exploration and learning of non-human-like reasoning paths, which ultimately improves performance. However, increasing the confidence threshold also introduces shorter lengths of chosen and rejected data, as shown in Figure 4. We argue that this is the reason for the performance decline of Qwen2-7B-SFT when the confidence threshold is greater. These results further demonstrate the necessity of optimizing the model by the intermediate reasoning path rather than by the entire process.
4.7 Analysis of Division Strategy
We use confidence scores to reconstruct the initial reasoning step , as well as the chosen and rejected data in the original Step-DPO-10k dataset, in order to compare human-like and non-human-like step divisions. Specifically, we truncate the initial reasoning step in the Step-DPO-10k dataset at the lowest confidence point, as in the dataset construction of CGPO. For the chosen and rejected data, we apply a truncation threshold at the lowest 2% confidence. The final results are reported in Table 5.
We find that even using $\pi_{0}$ to split the reasoning path on outputs generated by the strong model for training, the resulting performance still surpasses the baseline. However, it does not outperform training on the $\pi_{0}$ own generated and split data. This result suggests that the split of its own generated outputs by the weak model is more aligned with the model, thereby leading to improved training performance.
Table 5: Model performance with different annotation or split methods.
Model GSM8K MATH
MetaMath-Llama-8B   + CGPO 85.3 40.3
+ Step-DPO 84.8 40.2
+ Step-DPO-Confidence 84.4 40.2
Qwen2-7B-SFT   + CGPO 88.6 54.8
+ Step-DPO 88.4 54.6
+ Step-DPO-Confidence 88.1 54.5
5 Conclusion
In this paper, we propose CGPO, a DPO-based step-wise reasoning path preference optimization method. Our method emphasizes non-human-like reasoning paths in two key ways: 1) the training data does not rely on strong models; 2) the split for reasoning paths does not depend on predefined anchors. We show the effectiveness of CGPOin both mathematics and code domains, achieving significant improvements and generalization ability over approaches that use data generated by strong models. Furthermore, our experiments reveal that, within a certain range, intermediate reasoning path-focused optimization leads to greater performance gains than whole reasoning path optimization.
Limitations
Due to computational constraints, we are unable to conduct experiments on larger models (e.g.,70B) to explore the scalability with respect to model size. In addition, our experiments were limited to the code and math domains, primarily because fine-grained evaluation in the commonsense domain is challenging. However, we believe our approach remains applicable as long as a general fine-grained scoring model is available, since the proposed methodology is domain-agnostic.
Appendix A Appendix
A.1 Example of Confusion vs. First Error
We present two examples where the lowest confidence appears prior to the first error in Figure 5.
Figure 5: Examples where the lowest confidence appears before the first error.
A.2 Statistics of the Positional Relationship
Table 6: Statistics on the relative positions of first error step and lowest confidence point, where "Before" and "After" denote the first error before or after the model confusion point
Model Position
Before After
MetaMath-Llama-8B 22 78
Qwen2-7B-SFT 28 72
We randomly select 200 samples with incorrect results, half of which are from the MetaMath-Llama-8B generated CGPO training dataset and the other from the Qwen2-7B-SFT dataset. To identify the first erroneous step, we use GPT5-thinking with the prompt-
"Analyze the following text and identify the first location where an error occurs, reporting the exact substring/token, its position (line and character index), why it is an error, and why this is the first error (i.e., why all preceding text is error-free); if no error exists, reply ‘No error found’; Text:" [Q-A pair]
-to find the first error, and manually determine the positional relationship between the first error and the model’s lowest confidence point. We report the analysis results in Table 6.